# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 15 Trace maps
As $\mathcal{X}$ has semistable reduction, then $\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})$ is a semistable Galois representation and the cohomology $\mathrm{H}_{\mathrm{logcr}}^{1}\left(\mathcal{X}_{k}/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)$ of $\mathcal{X}_{k}$, without considering the horizontal divisors, is isomorphic to $D_{\mathrm{st}}\left(\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})\right)^{N=0}$, i.e., $D_{\mathrm{crys}}\left(\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})\right)$, by Fontaine’s $C_{\rm st}$-conjecture proved by Tsuji. We can identify $\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})$ as Galois representation with the rational $p$-adic Tate module $V_{p}(A)$ of $A=\mathrm{Pic}^{0}(X^{\rm alg}_{K}/K)$. Since $\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})$ is already semistable, possibly enlarging $K$ so that $A$ acquires semistable reduction, $V_{p}(A)$ is an extension of Galois representations
$0\longrightarrow V_{p}(G)\longrightarrow V_{p}(A)\longrightarrow Z\otimes\mathbb{Q}_{p}\longrightarrow 0,$
where $Z$ is a finite and free $\mathbb{Z}$-module, $G$ is a semiabelian scheme extension $0\to T\to G\to B\to 0$ of an abelian scheme $B$ by a torus $T$ over $\mathcal{O}_{K}$ and $Z\to G_{K}]$ is the $1$-motive giving Mumford’s unifomization of $A$. Let $q=p^{n}$ be the cardinality of the residue field of $K$. Then, $D_{\mathrm{crys}}\left(\mathrm{H}_{\mathrm{et}}^{1}(X_{\overline{K}},\mathbb{Q}_{p})\right)=D_{\mathrm{crys}}\bigl(V_{p}(G)\bigr)$ which is an extension of $D_{\mathrm{crys}}\bigl(V_{p}(B)\bigr)$ by $D_{\mathrm{crys}}\bigl(V_{p}(T)\bigr)$ and geometric Frobenius has eigenvalues which are Weil numbers of archimedean norm $q$ on $D_{\mathrm{crys}}\bigl(V_{p}(T)\bigr)$ and $q^{1/2}$ on $D_{\mathrm{crys}}\bigl(V_{p}(B)\bigr)$.
Since we have an exact sequence
$\mathrm{H}_{\mathrm{logcr}}^{1}\left(\mathcal{X}_{k}/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)\longrightarrow\mathrm{H}_{\mathrm{HK}}^{1}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)\longrightarrow\oplus_{j\in I}\mathrm{H}_{\mathrm{logcr}}^{0}\left(\mathcal{D}_{j,k}/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)(-1),$
geometric Frobenius acts as the identity on each $\mathrm{H}_{\mathrm{logcr}}^{0}\left(\mathcal{D}_{j,k}/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)\cong K_{0}$ and geometric Frobenius is $\Phi^{n}$, the claim follows. ∎
The reason behind the definition is the following Lemma.
@@
Assume that $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$. Then $P\left(\Phi\right)$ is isomorphism on $\mathrm{H}_{\mathrm{an},c}^{2d}$ and on $\mathrm{H}_{\mathrm{an},c}^{2d+1}$. In particular, if $\mathbb{I}$ is the unit object $(\mathcal{O},\mathfrak{O})$ then we get
$\mathrm{H}^{2d}_{\mathrm{dR},c}(U)_{L}\cong F^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right).$
Démonstration.
Condition (iii) of Definition 13 is equivalent to ask that $P(\Phi_{L})$ is an isomorphism on $(\mathrm{H}_{\mathrm{HK}}^{1,N=0})^{\vee}(-d-1)$. As the Hyodo-Kato cohomology with trivial coefficients also satisfies Poincaré duality, see [EY, Prop. 4.6], we have a perfect pairing $\mathrm{H}_{\mathrm{HK}}^{1}\otimes_{\mathbb{Q}_{p}}\mathrm{H}_{\mathrm{HK},c}^{2d-1}\rightarrow K_{0,L}\left(-d\right)$ and condition (iii) becomes equivalent to require that $P\left(\Phi\right)$ is an isomorphism on $\mathrm{H}_{\mathrm{HK},c}^{2d-1}(-1)/N\mathrm{H}_{\mathrm{HK},c}^{2d-1}(-1)$. It follows from Proposition 6.2 that $P\left(\Phi_{L}\right)$ is an isomorphism on $\mathrm{H}_{\mathrm{an},c}^{2d}$ (resp. $\mathrm{H}_{\mathrm{an},c}^{2d+1}$ ) if and only if it is an isomorphism on $\mathrm{H}_{\mathrm{abs},c}^{2d-1}$ (resp. $\mathrm{H}_{\mathrm{abs},c}^{2d}$). Here we write $\mathrm{H}_{\mathrm{abs},c}^{i}$ for the absolute cohomology $\mathrm{H}_{\mathrm{abs}}^{i}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0},\mathfrak{O}_{0}\left(-\mathfrak{D}_{0}\right)\right)_{L}$ of the structure sheaf and similarly we write $\mathrm{H}_{\mathrm{an},c}^{i}$ for the analytic cohomology with support $\mathrm{H}_{\mathrm{an}}^{i}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0},\mathfrak{O}^{\mathrm{conv}}\left(-\mathfrak{D}\right)\right)_{L}$ of the structure sheaf. By Proposition 6.2 this follows if $P\left(\Phi_{L}\right)$ is an isomorphism on the spaces $\mathrm{H}_{\mathrm{HK},c}^{2d-1}\left(-1\right)/N\mathrm{H}_{\mathrm{HK},c}^{2d-1}\left(-1\right)$ and $\mathrm{H}_{\mathrm{HK},c}^{2d}\simeq K_{0,L}\left(-d\right)$ (resp. $P\left(\Phi_{L}\right)$ is an isomorphism on $\mathrm{H}_{\mathrm{HK},c}^{2d}\left(-1\right)/N\mathrm{H}_{\mathrm{HK},c}^{2d}\left(-1\right)\simeq K_{0,L}\left(-\left(d+1\right)\right)$). Here we use that Hyodo-Kato cohomology is 0 in degrees >2d>2d and that we have an isomorphism $\mathrm{Tr}_{\mathrm{HK}}\colon\mathrm{H}_{\mathrm{HK},c}^{2d}\simeq K_{0,L}\left(-d\right)$ ; see the discussion in [EY, Prop. 4.6]. Since $P\left(\Phi\right)$ acting on $K_{0,L}\left(-t\right)$ is invertible iff $\alpha_{i}\neq p^{t}\zeta$ for every $i\in I$ and every $\zeta\in\mu_{f}$, these conditions amount to conditions (i), (ii) and (iii) of Definition 13. The first claim follows.
We prove the second claim. By Lemma 14.2 we have $F^{0}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=0$ so that $\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=F^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$ and the latter is identified with $\mathrm{H}_{\mathrm{logdR}}^{2d}\left(X,\mathcal{O}_{X}(-D)\right)_{L}$ which is isomorphic to $\mathrm{H}_{\mathrm{dR},c}^{2d}(U,\mathcal{O}_{U})_{L}$. ∎
We use this lemma to give the following definition. Let $\mathrm{Tr}_{U}\colon\mathrm{H}_{\mathrm{dR},c}^{2d}(U,\mathcal{O}_{U})\cong K$ be the coherent trace map of [LLZ, Thm. 4.2.7] and let $\mathrm{Tr}_{\mathrm{dR}}\colon\mathrm{H}_{\mathrm{logdR}}^{2d}\left(X,\mathcal{O}_{X}(-D)\right)\cong K$ be the induced trace map using that $\mathrm{H}_{\mathrm{dR},c}^{2d}(U,\mathcal{O}_{U})\cong\mathrm{H}_{\mathrm{logdR}}^{2d}\left(X,\mathcal{O}_{X}(-D)\right)$.
Definition 14.
Assume that $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$. We let
$\mathrm{Tr}_{P,r}\colon\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\longrightarrow K_{L}$
be the inverse of composite of the identifications $\mathrm{H}^{2d}_{\mathrm{dR},c}(U)_{L}\cong F^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$ and $\mathrm{Tr}_{U_{L}}$. More explicitly, following 14.2, it sends
$F^{1}\mathrm{H}^{2d+1}\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\ni(x,y)\mapsto\mathrm{Tr}_{\mathrm{dR}}\left(y-\gamma_{\mathrm{an},c}\left(P\left(\Phi\right)^{-1}x\right)\right)\in K_{.}$
Remark 15.3.
If $P\mid Q$ and $r\geq s$ and $\left(Q,s\right)$ is admissible, then the same holds true for $\left(P,r\right)$. In this case, $\mathrm{Tr}_{P,r}$ equals the composition of the morphism $(F^{1})\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\rightarrow(F^{1})\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},Q,s\right)\right)$ of Remark 14.1 followed by $\mathrm{Tr}_{Q,r}$.
We conclude this section with the following compatibility of trace maps. Let $m:=[K_{0}:\mathbb{Q}_{p}]$.
Lemma 15.4.
The following diagram commutes :
@@
${\mathrm{H}_{\mathrm{an}}^{2d}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\left(\mathfrak{S},\mathfrak{M}\right),\mathfrak{O}^{\mathrm{conv}}(-\mathfrak{D})\right)^{\Phi^{m}=p^{dm}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\wr}$${\mathrm{H}_{\mathrm{logdR}}^{2d}\left(X,\mathcal{O}_{X}(-D)\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${K}$${\mathrm{H}_{\mathrm{HK}}^{2d}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}(-\mathfrak{D}_{0})\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${K_{0}(-d),\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
where the horizontal isomorphisms are the trace maps $\mathrm{Tr}_{\rm dR}$ introduced above and $\mathrm{Tr}_{\rm HK}$ of [EY, Prop. 4.6] respectively, the left vertical map is defined by $Z=0$ and the map $K_{0}(-d)\to K$ is the natural inclusion $K_{0}\subset K$.
@@
Remark 15.5.
One may also allow $\mathcal{X}$ to be proper and of relative dimension $d$ and the same results hold. (By applying the above discussion to each $\mathcal{X}_{i}$’s in order to deduce that $P\left(\Phi_{L}\right)$ acting on $\mathrm{H}_{\mathrm{an},c}^{2d}$ is invertible when $\left(P,r\right)$ is admissible, that $F^{0}\mathrm{H}^{2d+1}\left(\mathbf{Syn}_{P,r}^{c}\right)=0$ and noticing that the existence of traces on each component imply that $\mathrm{H}^{2d}\left(\Gamma_{\mathrm{dR},L}^{c}\left(\mathbb{I}\right)\right)$ is a sum of copies of $K_{L}\left(d\right)$. The only difference is that $\mathrm{Tr}_{\mathrm{dR}}$ is a morphism which is not in general an isomorphism).
@@ 16.3 Cup products in syntomic cohomology
We go back to the notation of §16.1. Recall that, if $P_{1}\left(X\right),P_{2}\left(X\right)\in 1+XL\left[X\right]$ and write $P_{i}\left(X\right)=\prod\nolimits_{\alpha\in R_{i}}\left(1-\alpha X\right)$ where $\alpha\in\overline{L}$. Then
@@
We have that $\Cup_{P_{1},r_{1};P_{2},r_{2}}$ is a cup product of double complexes satisfying Leibniz rule ; see Definition 15. Moreover, setting $P_{3}:=P_{1}\ast P_{2}$ and $r_{3}:=r_{1}+r_{2}$, we have the following commutative diagram of complexes in which the horizontal arrows are the cup products of complexes satisfying the Leibniz induced by the respective $\Cup_{P_{1},r_{1};P_{2},r_{2}}$’s proceeding as explained after Definition 15 :
$\begin{array}[]{ccc}\mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}^{c\text{-}\infty}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}^{c\text{-}\infty}(\mathbb{E}_{3},P_{3},r_{3})\\ 1\otimes\tau_{\mathrm{rig,c}}\big\uparrow\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }&&\text{ \ \ \ \ \ \ }\big\uparrow\tau_{\mathrm{rig,c}}\\ \mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})\\ \tau_{\mathrm{rig}}\otimes 1\big\downarrow\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }&&\ \ \big\downarrow\wr\\ \mathbf{Syn}_{\mathrm{rig}}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3})\text{.}\end{array}$
Here $\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3}):=\mathbf{Syn}_{\mathrm{rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E}_{i}^{\dagger},\mathbb{E}_{i}^{\dagger},P_{i},r_{i})$ is the overconvergent version of $\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})$ defined as in Definition 6 using the $\mathfrak{E}_{\underline{\beta}}^{\dagger}$’s of §4.5. Moreover, we set $\mathbf{Syn}_{\mathrm{rig}}:=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}$, $\mathbf{Syn}_{\mathrm{rig}}^{c}:=\mathbf{Syn}_{\mathrm{rig},U^{o}\subset Y,\alpha_{k}}^{c}$ with $U=\mathcal{X}_{k}^{\mathrm{sm}}$, $U^{o}=U\backslash\mathcal{D}_{k}$ and $Y=\mathcal{X}_{k}$ and similarly $\tau_{\mathrm{rig}}=\tau_{\mathrm{rig},U\subset Y,\alpha_{k}}$ and $\tau_{\mathrm{rig,c}}$ are the morphisms of double complexes defined in §4.5 and in §6.3. We also have the variant given by replacing everywhere in the diagram $\mathbf{Syn}$ with $\mathbf{Syn}^{c\text{-}\infty}$.
@@
defines a cup product of double complexes satisfying Leibniz rule ; see [Stacks Project, (20.25.3.2)], where $p_{1}+p_{2}$ is denoted $p$, $p_{1}+q_{1}$ is denoted $n$ and $p_{2}+q_{2}$ is denoted $m$). In particular, we get cup products $\cup$ of complexes satisfying the Leibniz rule for these de Rham complexes $\Omega$. This means that, in order to check Definition 15 holds for the double complexes giving rise to $\mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})$ and $\mathbf{Syn}^{c\text{-}\infty}(\mathbb{E}_{2},P_{2},r_{2})$, it suffices to check the first relation, whose only non-trivial case is when $p_{1}=p_{2}=0$ in the table above and the required relation follows thanks to (10 ).
@@
$\mathrm{Cone}_{\underline{\beta}}\left(\mathfrak{E}_{i}^{{\dagger}}\right):=\mathrm{Cone}\left(-1:\mathrm{dR}\left(\mathfrak{E}_{i}^{?}\left(-\mathfrak{D}\right)_{\underline{\beta}}\right)\longrightarrow\mathrm{dR}\left(j_{\underline{\beta}\ast}\circ j_{\underline{\beta}}^{\ast}\left(\mathfrak{E}_{i}^{?}\left(-\mathfrak{D}\right)\right)_{\underline{\beta}}\right)\right)\left[-1\right].$
@@
by means of the formula $z\sqcup_{\underline{\beta}}\left(x,y\right):=\left(p_{1}^{\ast}\left(z\right)\wedge p_{2}^{\ast}\left(x\right),p_{1}^{\ast}\left(z\right)\wedge p_{2}^{\ast}\left(x\right)\right)$. For $R\Gamma_{\mathrm{rig}}\left((U^{o}\subset Y,\alpha_{k})/K,\mathcal{E}_{i}\right)$ and $R\Gamma_{\mathrm{rig},c}\left((U^{o}\subset Y,\alpha_{k})/K,\mathcal{E}_{i}\right)$ one argues similarly. As above, we can use ( 11) in order to get cup products of complexes satisfying the Leibniz rule and then get the required cup products of double complexes satisfying Leibniz rule from them. We remark that (12) for $\mathrm{conv}$ and $\dagger$ are compatible with respect to the restriction from the convergent to the overconvergent stuff. As the map $\tau_{\mathrm{rig}}$ is restriction to an open and $\tau_{\mathrm{rig,c}}$ is the projection of the cone onto the first component, the commutativity of the diagram follows. Finally, Corollary 6.4 gives the quasi-isomorphism $\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})\longrightarrow\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3})$. ∎

### v2

@@ 15 Trace maps
We have that $\mathrm{F}^{0}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\subset\mathrm{H}_{\mathrm{an},c}^{2d+1,P\left(\Phi_{L}\right)=0}$, hence it suffices to show that $\mathrm{H}_{\mathrm{an},c}^{2d+1,P\left(\Phi_{L}\right)=0}=0$. It follows from Proposition 6.2 $\left(3\right)$ and $\left(4\right)$ and the vanishing of $\mathrm{H}_{\mathrm{HK},c}^{2d+1}$ that we have
$\mathrm{H}_{\mathrm{an},c}^{2d+1,P\left(\Phi_{L}\right)=0}\simeq\mathrm{H}_{\mathrm{HK},c}^{2d}\left(-1\right)^{P\left(\Phi_{L}\right)=0}\simeq K_{0,L}\left(-d-1\right)^{P\left(\Phi_{L}\right)=0},$
which is zero since by assumption $P\left(\Phi_{L}\right)$ is an isomorphism on $K_{0,L}\left(-d-1\right)$. ∎
Let $\mathrm{Tr}_{U}\colon\mathrm{H}_{\mathrm{dR},c}^{2d}(U,\mathcal{O}_{U})\cong K(-d)$ be the coherent trace map defined in Equation $(4.2.8)$, Theorem 4.2.7 of [LLZ23], and define the de Rham trace map
$\mathrm{Tr}_{\mathrm{dR}}\colon\mathrm{H}_{\mathrm{logdR},c}^{2d}\cong K_{L}\left(-d\right)$
to be its composition with the isomorphism $\mathrm{H}_{\mathrm{logdR},c}^{2d}\cong{}\mathrm{H}_{\mathrm{dR},c}^{2d}(U,\mathcal{O}_{U})_{L}$. We can finally define the sought for finite polynomial trace map.
Definition 14.
Assume that $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$. Define
$\mathrm{Tr}_{P,r}\colon\mathrm{H}^{2d+1}\big(\mathbf{Syn}^{c}\big(\mathbb{I},P,r\big)\big)\longrightarrow K_{L}$
to be the composition of the natural isomorphism
$\mathrm{H}^{2d+1}\big(\mathbf{Syn}^{c}\big(\mathbb{I},P,r\big)\big)=\mathrm{F}^{1}\mathrm{H}^{2d+1}\big(\mathbf{Syn}^{c}\big(\mathbb{I},P,r\big)\big)$
of Lemma 15.2, the map $\partial$, and the de Rham trace map $\mathrm{Tr}_{\mathrm{dR}}$. Explicitly, after identifying $\mathrm{H}^{2d+1}\big(\mathbf{Syn}^{c}\big(\mathbb{I},P,r\big)\big)$ with $\mathrm{F}^{1}\mathrm{H}^{2d+1}\big(\mathbf{Syn}^{c}\big(\mathbb{I},P,r\big)\big)$, one has
$\mathrm{Tr}_{P,r}([x,y])=\mathrm{Tr}_{\mathrm{dR}}\left(y-P\left(\Phi_{L}\right)^{-1}\gamma_{\mathrm{an},c}\left(x\right)\right)$
[…]
Remark 15.3.
If $P$ divides $Q$ in $1+XL[X]$, $r\geqslant s$, and $\left(Q,s\right)$ is admissible, then $(P,r)$ is admissible, and $\mathrm{Tr}_{P,r}$ equals the composition of $\mathrm{Tr}_{Q,s}$ with the morphism $\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\longrightarrow\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},Q,s\right)\right)$ of Remark 14.1.
We conclude this section with the following compatibility of trace maps. Let $I^{\mathrm{an}}_{c}$ be the kernel of $\gamma^{\mathrm{an}}_{c}:\mathrm{H}^{2d}_{\mathrm{an},c}\stackrel{{\scriptstyle}}{{\longrightarrow}}\mathrm{H}^{2d}_{\mathrm{logdR},c}$. It is also the kernel of $\gamma_{c}^{\mathrm{HK}}:\mathrm{H}^{2d}_{\mathrm{an},c}\stackrel{{\scriptstyle}}{{\longrightarrow}}\mathrm{H}^{2d}_{\mathrm{HK},c}$ (defined by $Z=0$), as $\gamma^{\mathrm{an}}_{c}=\mathrm{comp}_{c}\circ{}\gamma^{\mathrm{HK}}_{c}$ and $\mathrm{comp}_{c}$ is injective.
@@
The following diagram commutes.
@@
${\mathrm{H}_{\mathrm{an}}^{2d}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\left(\mathfrak{S},\mathfrak{M}\right),\mathfrak{O}^{\mathrm{conv}}(-\mathfrak{D})\right)\big/I_{c}^{\mathrm{an}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma^{\mathrm{HK}}_{c}}\scriptstyle{\wr}$$\scriptstyle{\ \ \ \ \ \ \ \ \ \ \ \ \gamma^{\mathrm{an}}_{c}}$${\mathrm{H}_{\mathrm{logdR}}^{2d}\left(X,\mathcal{O}_{X}(-D)\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${K}$${\mathrm{H}_{\mathrm{HK}}^{2d}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}(-\mathfrak{D}_{0})\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${K_{0}(-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
Here the horizontal isomorphisms are the trace maps $\mathrm{Tr}_{\rm dR}$ introduced above and $\mathrm{Tr}_{\rm HK}$ of [EY20, Prop. 4.6] respectively, and the map $K_{0}(-d)\longrightarrow{}K$ is the inclusion.
@@ 16.3 Cup products in syntomic cohomology
We go back to the notation of §16.1. Recall that, if $P_{1}\left(X\right),P_{2}\left(X\right)\in 1+XL\left[X\right]$ and write $P_{i}\left(X\right)=\prod\nolimits_{\alpha\in R_{i}}\left(1-\alpha X\right)$ where $\alpha\in\overline{L}$, then
@@
We have that $\Cup_{P_{1},r_{1};P_{2},r_{2}}$ is a cup product of double complexes satisfying Leibniz rule ; see Definition 15. Moreover, setting $P_{3}:=P_{1}\ast P_{2}$ and $r_{3}:=r_{1}+r_{2}$, we have the following commutative diagram of complexes in which the horizontal arrows are the cup products of complexes satisfying the Leibniz induced by the respective $\Cup_{P_{1},r_{1};P_{2},r_{2}}$’s (via the associated $\widetilde{\Cup}_{P_{1},r_{1};P_{2},r_{2}}^{p_{1},q_{1};p_{2},q_{2}}:=\left(-1\right)^{q_{1}p_{2}}\Cup_{P_{1},r_{1};P_{2},r_{2}}^{p_{1},q_{1};p_{2},q_{2}}$, see the discusion before Proposition 16.1) and where the morphism
$\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})\longrightarrow\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3})$
is a quasi-isomorphism :
$\begin{array}[]{ccc}\mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}^{c-\infty}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}^{c-\infty}(\mathbb{E}_{3},P_{3},r_{3})\\ 1\otimes\tau_{\mathrm{rig,c}}\uparrow\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }&&\text{ \ \ \ \ \ \ }\uparrow\tau_{\mathrm{rig,c}}\\ \mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})\\ \tau_{\mathrm{rig}}\otimes 1\downarrow\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }&&\downarrow\\ \mathbf{Syn}_{\mathrm{rig}}(\mathbb{E}_{1},P_{1},r_{1})\otimes_{K_{0},L}\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{2},P_{2},r_{2})&\longrightarrow&\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3})\text{.}\end{array}$
Here $\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3}^{\dagger},P_{3},r_{3}):=\mathbf{Syn}_{\mathrm{rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathbb{E}_{i}^{\dagger},P_{i},r_{i})$ is an overconvergent version of $\mathbf{Syn}_{\mathrm{rig}}^{c}(\mathbb{E}_{3},P_{3},r_{3})$ (obtained considering the $\mathfrak{E}_{\underline{\beta}}^{\dagger}$’s of §4.5 and replacing the $\mathfrak{E}_{\underline{\beta}}^{\mathrm{conv}}$’s with them in Definition 6) and we have $\mathbf{Syn}_{\mathrm{rig}}:=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}$, $\mathbf{Syn}_{\mathrm{rig}}^{c}:=\mathbf{Syn}_{\mathrm{rig},U^{o}\subset Y,\alpha_{k}}^{c}$ with $U=\mathcal{X}_{k}^{\mathrm{sm}}$, $U^{o}=U\backslash\mathcal{D}_{k}$ and $Y=\mathcal{X}_{k}$. Furthermore, $\tau_{\mathrm{rig}}=\tau_{\mathrm{rig},U\subset Y,\alpha_{k}}$ and $\tau_{\mathrm{rig,c}}$ are the morphisms of double complexes defined in §4.5 and in §6.3. We can also replace everywhere in the diagram $\mathbf{Syn}$ with $\mathbf{Syn}^{c\text{-}\infty}$.
@@
defines a cup product of double complexes satisfying Leibniz rule ; see [Stacks Project, (20.25.3.2)], where $p_{1}+p_{2}$ is denoted $p$, $p_{1}+q_{1}$ is denoted $n$ and $p_{2}+q_{2}$ is denoted $m$). In particular, we get cup products $\cup$ of complexes satisfying the Leibniz rule for these de Rham complexes $\Omega$. This means that, in order to check Definition 15 holds for the double complexes giving rise to $\mathbf{Syn}(\mathbb{E}_{1},P_{1},r_{1})$ and $\mathbf{Syn}^{c-\infty}(\mathbb{E}_{2},P_{2},r_{2})$, it suffices to check the first relation, whose only non-trivial case is when $p_{1}=p_{2}=0$ in the table above and the required relation follows thanks to (11).
@@
$\mathrm{Cone}_{\underline{\beta}}\left(\mathfrak{E}_{i}^{{\dagger}}\right):=\mathrm{Cone}\left(-1:\mathrm{dR}\left(\mathfrak{E}_{i}^{?}\left(-\mathfrak{D}\right)_{\underline{\beta}}\right)\longrightarrow\mathrm{dR}\left(j_{\underline{\beta}\ast}\circ j_{\underline{\beta}}^{\ast}\left(\mathfrak{E}_{i}^{?}\left(-\mathfrak{D}\right)\right)_{\underline{\beta}}\right)\right)\left[-1\right]\text{.}$
@@
by means of the formula $z\sqcup_{\underline{\beta}}\left(x,y\right):=\left(p_{1}^{\ast}\left(z\right)\wedge p_{2}^{\ast}\left(x\right),p_{1}^{\ast}\left(z\right)\wedge p_{2}^{\ast}\left(x\right)\right)$. For $R\Gamma_{\mathrm{rig}}\left((U^{o}\subset Y,\alpha_{k})/K,\mathcal{E}_{i}\right)$ and $R\Gamma_{\mathrm{rig},c}\left((U^{o}\subset Y,\alpha_{k})/K,\mathcal{E}_{i}\right)$ one argues similarly. As above, we can use (12) in order to get cup products of complexes satisfying the Leibniz rule and then get the required cup products of double complexes satisfying Leibniz rule from them. We remark that (13) for $\mathrm{conv}$ and $\dagger$ are compatible with respect to the restriction from the convergent to the overconvergent stuff. As the map $\tau_{\mathrm{rig}}$ is restriction to an open and $\tau_{\mathrm{rig,c}}$ is the projection of the cone onto the first component, the commutativity of the diagram follows. Finally, Corollary 6.4 proves that we have the claimed quasi-isomorphism. ∎